2009-Q3

Following the links below, you can read the detailed reports submitted for the third quarter 2009. 2009 Q3-1 -- OpEng / Andres Gonzalez: 2009 Q3-1 -- DIMM / Ariel Sanchez 2009 Q3-1 -- 2.2m / Francisco Labrana 2009 Q3-1 -- FEROS / Monica Castillo 2009 Q3-1 -- WFI / Pablo Arias 2009 Q3-1 -- 3.6m / Eduardo Matamoros 2009 Q3-1 -- HARPS / Gino Leon 2009 Q3-1 -- NTT / Leonardo Gallegos 2009 Q3-1 -- EFOSC2 / Francisco Caceres 2009 Q3-1 -- SOFI / Manuel Pizarro

2009 Q3-1 -- OpEng

OPERATIONS ENGINEER Full Quarterly Report July – August - September 2009 TABLE OF CONTENTS 1. HIGHLIGHTS……………………………………………………………………...….3. 2. DETAILED REPORTS…………………..………3. 3. PROJECT FOR NEXT PERIORD…………………………………………….……………………3. 4. CONCERNS...……………………………………………………………….4. 1. HIGHLIGHTS Coordinate the removal and installation of the three telescopes consoles from the RITZ to the new control room (NOB). Modifications osf Startup according new control configuration from 3 telescopes. First visitor astronomer introduction according new model la Silla 2010+ 2. DETAILED REPORTS 2.1. Operation Engineer tasks : TIOS day-operations training: - Startup Telescopes (NTT, 3.6, 2.2) - Day activities: Tasks coordinated with other teams, CCD test, and Instruments calibrations - Instruments setup, trouble-shotting 2.1.1 Sciops Coordinator 2.1.2 Templates problems and improvement 2.2 Operations Engineer Project 2.2 AutRep: Development of different scripts for Telescopes and instruments to check and monitor several parameters to avoid problems on the future. See http://regle.ls.eso.org/autrep2/pages/ 2.2.2 Training for Visitor astronomer introduction with P2PP 3. NEXT PERIOD PROJECT Working on a new general panel to monitor the Telescopes and Instruments status, all at the same time. Starting with new activities to technical support for visitor astronomer (P2PP, GASGANO, OT) Autrep: continue on script development to monitor and preventive maintenance of telescopes and instruments. I will work in EFOSC2 script to Zero Points 4. CONCERNS According to the La Silla 2010+ needs and me news duties and responsibilities, I would suggest to have and astronomer “on call” to clarify any astronomical question ___________________________________oOo______________________________________

2009 Q3-1 -- DIMM

La Silla Observatory: Quarterly Report Instrument Dimm Name: Ariel Sanchez Period: Q3 2009 (July - September 2009) Date: Q3 September 2009 ------------------------------------------------------------------------------ STATUS: - Personnel: IOT: Instrument TIO: Ariel Sanchez Instrument Friend: Ariel Sanchez Software : Alex Segovia - General: During the ESO Weather Watch (PWV) Campaign : (http://www.eso.org/gen-fac/pubs/astclim/forecast/meteo/update2000/) A number of issues were detected with the data not being accessible (it seems not archived). The requirement is to investigate why the data is not accessible, and why it appears there is no correct archiving, as this is important for the Campaign (which is returning in October). This problem is assigned to Guillermo Valdes. Meteomonitor web page was not refreshing data. - Instrument Report Data is not accesible because is still not generated. The meteo.log file will be generated and transfered during the next weeks, and is for sure will be available for the next PWV campaing. Florian Kerber added to cc (responsible for PWV) The meteomonitor web page was not refreshing data because meteo.last file was not generated continuously by asmmetControl process (running on wlsops3 development workstation). It was caused by an electrical energy cut, then Vaisala meteo station lost telnet communication with workstation. To solved this problem, it was necessary restart asmmetControl process and then it recover telnet communication with Vaisala meteo station. * Open or assigned LPRS (ordered according to decreasing priority): 010047 : Meteo Station operational issues. Assigned 010053 : Meteomonitor webpage no updating. Closed * Other modified LPRS during Q3: 010072 : new camera looking at NTT is not working. Closed ------------------------------------------------------------------------------ AREAS OF CONCERN: ------------------------------------------------------------------------------ PROJECTS FOR NEXT PERIOD: Waiting results of the Up-grade of ParanalŽs dimm, this is scheduled for October 2009, If successful, will be the upgrade of La SillaŽs dimm. ------------------------------------------------------------------------------ USAGE OF TECHNICAL NIGHTS: Under planing by software ------------------------------------------------------------------------------ USAGE OF CALIBRATION PLAN NIGHTS (e.g. RESULTS, TRENDS, HEALTH CHECKS): n/a ------------------------------------------------------------------------------ ---oOo---

2009 Q3-1 -- 2.2m

La Silla Observatory: Quarterly Report Telescope 2p2 Name: Francisco Labrana Period: Q3 2009 (July-Sept) Date 28 Sept 2009 ----------------------------------------------------------------------------------------------------------------------------- STATUS: Telescope Report -Open or assigned LPRS -Closed LPRS -LPRS-010004 : VME reset Create Date: 07-07-2009 Detailed Description : While observing with FEROS the TCS VME did a reset for unknown reason, and the panels of the TCS, TCS Status and TCS Setup were completely white, then it was necessary a manual reset of the VME. Solution: This is re-occurrence of LPRS-009335 "TCS LCU self reboots". Time of occurrence is 05:57:49 UT. Like always this happened after focus preset. -LPRS-010015 : It is not possible to open the dome Create Date : 12-07-2009 Detailed Description : It was not possible to open the dome remotely. I went to the telescope and tried to open it locally but it didn't work. After that, I saw the chain broken. Solution: The slit mechanism was repaired by Ramon Leyton and Ignacio Lopez - LPRS-010017 : Control card of motor #4 failure Create Date : 12-07-2009 Detailed Description : Motor #4 was disabled until the card is repaired. Dome working with 3 motors for the time being. Solution: The board was changed and this fixed the problem. Training - CCD Astronomy by Blair Conn - HARPS and High Resolution Spectroscopy by Gaspare Lo Curto - Templates Course by Paul Le Saux

2009 Q3-1 -- FEROS

La Silla Observatory: Quarterly Report Instrument Feros Name: Monica Castillo Period: Q3 2009 (July-Sept) Date 28 Sept 2009 ----------------------------------------------------------------------------------------------------------------------------- STATUS: Instrument Report -Open or assigned LPRS --010058 : Feros Calibration unit Create Date : 03-09-2009 Detailed Description : Oliver Schuetz wrote Today the results of the FEROS wavelength calibration gave quite bad rms results, requiring an inspection of the calibration unit. Indeed, the calibration light beam did not hit the fibre entrances, but was about 1-2 cm lower. I tried to fix the ThAr lamp again, with a strange construction (which partly was already in use). The beam was then centered as good as possible on the fiber entrances. A lot of fine-tuning in the RMS encoder units - with the software interface - kept us busy all the afternoon. When I had to stop due to begin of the night, the obj fiber had very good rms quality, but the sky fiber showed still a bad S-shape. There is a chance that this S-shape could disappear in the next days and calibrations may be good. Status: In coming weeks and months, the calibration unit will likely disalign itself again. Monitoring needed. I will prepare a webpage as a guide how and what to check in the FEROS QC plots. A new calibration unit should soon be ready ? -Closed LPRS --010006 : Communication problems between Cam, Tcs, Fiera. Create Date: 08-07-2009 Detailed Description: After the Feros focus and when OBs started, all the system failed. All the system was restarted several times (TCS and FEROS rebooted because some zombies processes were found, Environments restarted, Shutdown and StartUp of Instrument and Telescope) but it did not solve the problem. M. Pavez was called. The following error message was found: Wferos Boss FeoControl BossINTERFACE_DCS. BossERR_GENERAL: No information about the failure of 'FIERA'. Solution: But at the next day, in the morning, all worked fine, and officially nothing was repaired, moved, fixed, etc. Network problems because Grond works ? Hard to investigate or reproduce it. Problem Closed. --010039 : FEROS Linearity Test OB failing and freezing BOB. Create Date: 04-08-2009 Detailed Description: The FEROS Linearity test is consistently killing the BOB when it reaches the end of the OB. It completes ok and calculates the linearity, but then it hangs and BOB becomes frozen and won't refresh at all. Possibly it is trying to raise a window and it can't. This seems to be the main problem with the old system trying to work on the new system. The only solution is to kill BOB and you may need to restart OT as well. Solution: Problem is solved, the problem happened while the template try to send the email. Thanks to A. Segovia for his support. --010041 : Process tifCA not active Create Date: 06-08-2009 Detailed Description: When the OB started on Feros, the following error was got: "process tifCS not active". The following errors messages were found on LogMonitor: qsemuThread.c : process not in environment REASON > process tifCA not active Cannot get information on process Could not send command "SETFIB OBJFIB" to w2p2tcs/tifCA. Solution: There is no other occurrence of this error during the last year and it is indeed a little odd. Apparently it was caused by several BOB instances running simultaneously (PIDs 19613, 25725, 30479). --010065 : Feros fiera shutter failed Create Date: 07-09-2009 Detailed Description: During a sequence Feros daily maintenance OB, the Fiera shutter failed sometimes. PSI was called, and he solved the problem. Solution: The shutter driver box appeared to be dead. No lights or actions from external controls. I replaced it with a spare one and the shutter control was restored. Yesterday the failing unit was inspected by J. Vilaza and he found that the mains power cable was broken inside the insulation at the place where it enters the box. It was repaired and the original shutter driver box re-installed and tested Ok. ___________________________________________________________________________ Training - CCD Astronomy by Blair Conn - HARPS and High Resolution Spectroscopy by Gaspare Lo Curto - Templates Course by Paul Le Saux

2009 Q3-1 -- WFI

Pablo Arias La Silla Observatory: Quarterly Report Instrument WFI Name: Pablo Arias Period: Q3 2009 (Jul -- Sep 2009) Date: September 27, 2009 ------------------------------------------------------------------------------ STATUS: - Personnel: I am still been attending to English classess at La Silla Observatory and ESO office in La Serena. Currently, I've been operating the 3.6 and 2.2 telescopes. I had my first training at NTT Telescope. Continuing with the courses on basic astronomical techniques, I've done the following: “CCD Astronmy” by B.Conn, “Harps and High Resolution Spectorscopy” by G.LoCurto and “Template Course” by P.LeSaux. IOT: Instrument TIO: Pablo Arias System Engineer: A.Pizarro/A.Macchino Instrument Friend: Pablo Arias - General: The behaviour of WFI is stable and there are not many problems like in the last report. Actually, on this period the majority problems were detected on the WFI images transfering. The Instrumentation and Software teams have been working on this problem. - Instrument Report * Problems with the transfers of the WFI's images happened during the readout and transfering (So it never completed the readout and the WFI's panel froze). It was not possible to abort the exposure. There are new occurrences of the same problem on the following tickets: LPRS-010040 and LPRS-010018. Monitoring. * Some errors were found with the flat field time (the calculated exposure time). A lamp was replaced from 650W to 1000W and the exposure times now came back to normal. Closed. * Trends of instrument health-check parameters WFI: flat, dome, bias, twilight flats; In general on this period, they have been keeping stable. * Open or assigned LPRS (ordered according to decreasing priority): -- LPRS-010059: “WFI image transfering not completed”; During the readout and transfering of the WFI's images the system froze (it never completed the last redout). It was not possible to abort the exposure. There are new occurrences of the same problem LPRS-010040 and LPRS-010018. The support was called but there was no way to unblock the system, turn OFF and ON the FIERA electronics did not help and only a SPARC reboot solved the problem. The Instrumentation team have tried to associate the date in which the errors appear but it did not match with interventions in WFI. During 2006, there was a LPRS-8089 and seems to be the same problem. At least DMA errors can be seen at syslog of SPARC in same way: "timeout expired waiting on DMA ready" and "DMA timeout expired". Checking shared memory (ipcs -m) that there was three unused memory block occupied with size of WFI image. They were removed (ipcrm -m <ID>. There was not bad shared memory blocks. The driver of the FIERA was re-installed and finally sparc changed, and after this to failures were not any more reported. It will keep on under investigation. -- LPRS-010022: “All WFI programs on w2p2off down”; Since the removal to the NOB, all Midas-based routines for WFI on w2p2off were down (Reduction of daily health check, Reduction of weekly health check and WFI Amag (derivation of photometric zero points)). The first 2 processes run nominally but produce completely empty output, while the Amag could not be started. All w2p2off Midas software needed to be adapted. The problems were fixed, setting new display on script to open midas in w2p2off. Daily and Weekly health check reduction was running ok. The display was setting also to WFI Amag and now is functioning. The daily and weekly health are working fine, and the newly implemented new_daily_health_check is working too. The printing is running ok in duplex and the right format. Under Monitoring. * Other modified LPRS during Q3: -- LPRS-010031: “Flat field time incorrect”; There was a problem with the calculated exposure time for WFI flatfields. When the software calculates the time which was needed to reach the desired flux level (usually 20000 counts in case of WFI, as given in the OBs), errors of a factor of 2 in the exposure time happened in approx. 20% of all cases. This problem is fully at random. Due to that a lamp was replaced from 650W to 1000W and exposure time now back to normal (as was being used before). Closed. -- LPRS-010042: “WFI CCD page not updating”; The online reporting of the CCD bias, ron, gain and flat values doesn't seem to have been updating. Now, The web page is http://www.eso.org/sci/facilities/lasilla/sciops/CCDs/WFI/DAILY/DHC/RON.gif. Closed. ------------------------------------------------------------------------------ AREAS OF CONCERN: ------------------------------------------------------------------------------ PROJECTS FOR NEXT PERIOD: ------------------------------------------------------------------------------ USAGE OF TECHNICAL NIGHTS: ------------------------------------------------------------------------------ USAGE OF CALIBRATION PLAN NIGHTS (e.g. RESULTS, TRENDS, HEALTH CHECKS): ------------------------------------------------------------------------------ ---oOo---

2009 Q3-1 -- 3.6m

La Silla Observatory: Quarterly Report Instrument 3p6 Telescope Name: Eduardo Matamoros Period: Q3 2009 (July -- September 2009) Date: September 29, 2009 ------------------------------------------------------------------------------ STATUS: - Personnel: A TIO (Cristian Romero) from PAO came to La Silla for three months, he started During September and he is working during daytime for medical recomendation. IOT: Instrument Scientists: Gaspare Lo Curto System Engineer: Jorge Vilaza Instrument Friend: Eduardo Matamoros - General: Problems in this period with the dome, related with encoder of the dome position The system was aligned and the zero position retrieved. After this intervention no other ocurrence. * Open or assigned LPRS (ordered according to decreasing priority): • Other modified LPRS during Q3: LPRS-010011 Hydraulic system down Closed Mechanical During a presetting the Hydraulic system down. To recover the system was necessary go to the 3th floor and run the procedure to put online. “7/15/2009 1:31:20 PM bahumada Fasteners of fixation of electric connectors of the pressure regulator of flow were adjusted side East number 1 at the 6 that were loose, orings of pressure regulator of flow West number 5 they were also changad” LPRS-010028 Computer Hardware Assigned TCS Panel with problems TCS panel has problems in the most important area of the panel Telescope Position (coordinates, sidereal time, HA, airmass, altitude, etc) isn't clear enough. Please try to fix it Marcus Pavez fixed problems, Now is OK. LPRS-010032 Mechanical Closed During a preset the Dome don’t stop at the new position. 7/27/2009 11:09:31 AM 3p6 During a preset dome don't stop in the new preset, log monitor give the following message: Alarm high activated for wheel 16 (24.71 <= 24.0 ) for load cell Alarm high activated for wheel 24 (25.08 <= 24.0 ) for load cell UT: 08:01:54 UT: 08:31:55 “8/11/2009 1:09:47 PM gihle There is no indication that the problem has a mechanical cause.” LPRS-010037 Software Closed TCS software failed due to full HD 8/4/2009 8:25:06 AM 3p6 Due to full HD in tcs machine, the DB snapshots files were corrupted. So, autoguider did not work. Then some devices started to fail, specially CCDimg camera. Several test was done by Alex Segovia and Andres Gonzalez. They find the solution after hours of fighting. “8/4/2009 4:47:17 PM asegovia The w3p6tcs ws /diska was 100% of it's capacity producing a general tcs and cam environments failure. Both, tcs and cam environments was not able to have operational their OLDB due to corrupted snapshots files, so a failure of the scanning systems to the LCUs related was produced. By cleaning the /diska (moving files to /diskb), was possible to recreate again the snapshots files and have tcs and cam environments working fine. In the meantime, we had problems to initialize the Adapter and Dec axis, so, for that reason it's took more time. A periodically disks space check is mandatory. Thanks a lot to Andrés Gonzalez, Gino León and Ismo for their support. LPRS-010061 Software Closed The 3p6 dome stuck 9/5/2009 11:14:44 AM 3p6 At 00:03 UT During the presetting a problem was detected with the dome because it never got the target position when we turned on the neon lamps We saw that it didn't move then G.LoCurto went to the telescope and moved it locally. We tried to move it several times from TCS and BOB but We got a timeout on the dome presetting. IKastinen solved the problem. “9/6/2009 2:48:31 PM ikastine Gaspare said that dome was moving really slow and was not reaching the dome reference point, and even was going away from it. Dome recovery buttons at TCS panel was tried to solve the problem. Seems that dome software was in strange state and after this it was not possible to move it via software, it could be moved only manually. In this point I came to control room. Pablo showed that there were errors in LCU prompt form domePresser function. In this point I re-initialized the dome software module, but this did not help. Then I pressioned the "dome recovery, No move" button and this released the dome control and it started to move reference point. I have not seen this kind dome going away from reference point behaviour before. Please, reopen if happens again.” LPRS-010066 Electronic Closed The dome moves slowly 9/8/2009 9:33:29 AM parias At 02:00 UT RA 205825.0, Dec 105021.0, HA 010816.0 During the observations (4/4 exposures) the star disappeared so I checked the dome status panel and I detected that the value of the dome New Position wasn't the same with the Dome Position (deg). The OB was aborted and the telescope moved to zenith and I called J.Vilaza. He went to the telescope to check the problem and He detected that there is some problem with the encoder. “9/10/2009 1:46:35 PM jvalenzu When turning the dome, no activity was displayed on the dome position LED's and software. We proceeded to dismount the encoder for testing it without load and actually it was working fine. By proceeding to remove the gear reduction, we realized that axis was not correctly coupled on the traction mechanism which is receiving the dome rotation. The system was aligned and the zero position retrieved. No problems reported. under monitoring some days. Thanks to B. Ahumada and O. Varas. ------------------------------------------------------------------------------ AREAS OF CONCERN: ------------------------------------------------------------------------------ PROJECTS FOR NEXT PERIOD: No meeting during last three months ------------------------------------------------------------------------------ USAGE OF TECHNICAL NIGHTS: No information about OIT ------------------------------------------------------------------------------ USAGE OF CALIBRATION PLAN NIGHTS (e.g. RESULTS, TRENDS, HEALTH CHECKS): No ------------------------------------------------------------------------------ ---oOo---

2009 Q3-1 -- HARPS

La Silla Observatory: Quarterly Report Instrument: HARPS Name: Gino Leon Period: Q3 (July -- September 2009) Date: 2009-09-30 ------------------------------------------------------------------------------ STATUS: IOT: Instrument Scientist: Gaspare LoCurto Instrument TIO: Gino Leon System Engineer: Peter Sinclaire Instrument Friend: Andres Gonzalez - Instrument Report * Trends of instrument health-check parameters Period Averages CCD1 temp: 148.156 [K] CCD2 temp: 148.216 [K] VV pressure : 5.17E-3 [mb] CCD temp averages [K] (CCD1) July : 148.15 August: 148.15 September: 148.15 CCD2 July : 148.15 August: 148.15 September:148.23 Vacuum Vessel pressure: [mb] July : 6E-3 August: 5E-3 September:5E-3 Harps Vacuum Pressure and ccd's pressure A raise of pressure and change of temperatures in CCD's happened in Harps technical time, on Sept. 5, maybe due to opening of vacuum vessel, to test Fabry Perot interferometer. These test were performed by Gaspare, but went back to normal quickly. A increase of CCD pressure happened from 1.15 nb, to 2 nb, changing again to 1.3 nb in a few hours. Open or Assigned problems * LPRS-010023 : Night log data did not appear. One night observing with p2pp, data did not get Remedy nightlog system. A crontab was installed to check system-health and automatic-restart. This problem should be closed, because did not happened again. Closed problems in this period * LPRS-010020 : "scratch" appearing in CFA fibre head plate. Apparently, there is some contamination over Harps fibre head plate, like a hair, or something like. Did appear on AG image, but did not bother, because AG worked anyway without problems. Soon it will be cleaned, when instrumentation people dismounted all to do manteinance. * LPRS-010021 : HARPS Custom pipeline stuck. Harps pipeline did not work, perharps due to afternoon EGGS corrupted calibration file. Next morning, it was copied new calibration files on a directory, and it was re-reduced all over again. * LPRS-010027 : WARMUP time of ThAr2 lamp to be increased. It was a petition to increase Th-Ar2 warm-up time, from 120s to 900s, to avoid a non-stability of new brand's lamps. * LPRS-010034 : OT crashes frequently because connection to database closes. It was a communication problem with database, due to firewall 's protection. Since did not appear more ocurrences, it was closed. * LPRS-010073 : harps vacuum sensor channel 2 give Error. A power cycle fixes the problem. * LPRS-010075 : thorium lamp 5 light intensity increassing during the night. It was known that happens when these lamps are closing to his dead. That happens probably for a plenty of calibrations done in Fabry-Perot testings. It was replaced and adjusted. ------------------------------------------------------------------------------ PROJECTS FOR NEXT PERIOD: - A tool for calculate mean values, dispersion and more, in autrep plots. - Measure Harps efficiency and comparing with previous data. ------------------------------------------------------------------------------ ---oOo---

2009 Q3-1 -- NTT

Ntt_2009_Q3 La Silla Observatory: Quarterly Report Instrument : NTT Name: Leonardo Gallegos Period: Q3 2009 (July 2009 -- September 2009) Date: 30 September 2009. ------------------------------------------------------------------------------ STATUS: - Personnel: Training in HARPS, Templates,CCD´s, Linux, English. IOT: Instrument TIO: Leonardo Gallegos Instrument Friend: Leonardo Gallegos * General: NTT hasn`t have relevant problem during this period, mainly the LPRS genereted has been associated to the autoguider system but it’s been solved promptly. Another significant situation was the GPS time server failure described in LPRS- 01030 below. Active optics has given very good results after to be applied, it has helped to have a good observations. * Instrument Report <> * Trends of instrument health-check parameters <> * Open or assigned LPRS (ordered according to decreasing priority): --010045 "Autoguider Camara Side A turned off several times" Autoguider side A turned off several times during the observation. Responsable A.Pizarro August 11,2009. No news has been reported. * Other modified LPRS during Q2: * Closed LPRS during Q3: -- 009960: “One Altitude encoder lamp is burned. The problem was detected during the telescope StartUp, because the Status "ALT encoder opMode" changed from "ALL 4 HEADS" to "BOARD 1 H1H2" in red color. Jorge Vilaza checked the lamps and found the H3 lamp burned.”Responsable Jorge Vilaza, April 18, 2009. LPRS closed, solution: All the 4 encoder heads in altitude have been refurbished with new lamps and properly aligned.Encoder calibration procedure done by IKA. In addition, 4 burned heads got new lamps and corresponding alignment. So, at present we have the altitude axis fully functional. no new occurrence has been reported since the intervention in july 6,2009. -- 010030 "Timeserver down due to UPS failed" Due to a failure in the small UPS connected to the GPS , Time server system and NTT time distribution , time server went down. Responsable Aldo Pizarro, July 25,2009. LPRS closed, solution: After to connect the system to the stabilized 220v line, many time was taken by the GPS to synchronize with satellite and several iteration were done after GPS was stabilized; that was the reason for the down time.Antenna was left connected for the night but we will keep this under monitoring. No new occurrency of this problem, and time it is syncronized. Aldo Pizarro August 8,2009. -- -- 009991: “The autoguider camera B is changing of gain level for the same guide star, is randomly jumping between 2 different gain level” Responsable Aldo Pizarro(june 11,2009). LPRS closed, solution: AGB ARC controller was replaced by a spare unit fixing the problem. no new occurrence has been reported since the intervention in july 14,2009. -- 010056 "LCU rotator side B with problem" Responsable. J.Vilaza Strange status in amplificador of co-rotator, it produced a interlock in the rotator. LPRS closed,solution:This problem was solved turning off the amplifier. J.Vilaza September 9,2009. LPRS-010062 "Amplifier of co-rotator side A with ploblem" Responsable J.Vilaza September 7,2009. During the first preset in the night rotator side A went down due to problem with co-rotator. LPRS closed, solution: all the electronic system was checked and it was ok, then just was neccesary to re-init the amplifier of co-rotator. J.Vilaza September 7,2009. No new occurrence. -- 010043 "AGA NTCCD continuos readout mode turned off several times during observation" It happened at 23:35 UT August 09,2009. Also to give more information the templates used were: SOFI_spec_obs_AutoNodOnSlit. SOFI_spec_cal_Arcs. Responsable Aldo Pizarro, August 9,2009. LPRS closed, solution: Changed flag to define if CCD must always automatically restarted with a full chip exposure when AG is not active (i.e. now autoRestart = 1). Rebuild tcs environment. Tested at night and no side effects reported.A.Segovia, August 17, 2009. -- 010063 "Scanlinks CCD autoguider side A down" Responsable P.Sinclair It is not clear the origin of this problem, but the indeed there was a failure at the scan links level from the lt5aga. Reassigned to software LPRS closed, solution: It was neccesary to re-init the scanlinks system. I.Kastine September 8,2009. -- 009984: “During spectroscopic observations using SofI, the target jumped from the chosen position on the slit in several occasions to bigger values of Y, about 20 pixels. In two of this occasions the telescope was pointing between 7 and 8 degrees from the zenith. It happened with and without the autoguider” Responsable Mario Alfaro, June 3, 2009. During a technical time were done some test about this apparent movement. Was found that it is not correlation between the position of the telescope and this wrong movement. This problem requires more testing as it is not yet clear what is exactly happening and under which conditions Also this problem could be associated to the altitude encoder lamp. LPRS closed, solution: No input has been received.Gerardo Ihle, September 15,2009. -- 009975: “M3 initialisation fails because the altitude is too far from zenith. Responsable Ismo Kastine. May 18, 2009 Was opened and Installed branch ntt 8.19.1.1. Tested, installed and archived ntt 8.19.1.1 ticket on hold and assigned to ismo for rollback once the encoder is fixed. LPRS closed, solution: The encoder was fixed, and change wqas roll back during NOB transfer. Ismo Kastine, July 29,2009. no new occurrence has been reported since the intervention. -- 010050 "The RTD of side B froze" Responsable J.Urrutia August 8,2009 This is apparently the same problem that affected 3p6. ASE will install the same patch and test it when SOFI is used. More thorough tests could be set up for the tech night on Sep 4th. LPRS closed, solution: This shift TCCD (side B, EFOSC observations, in this case) worked without any inconvenient, no new occurrences reported, and no comments from SCIOPS either. I observed it's behavior last night with TIO and it was satisfactory. So I decided not apply the same solution that I applied to 3p6 ccdIMG, until collect more details of the problem and be sure. Please Ismo check it by me this shift.A Segovia September 9,2009. Checked this shift with TIO, no new occurrences. --009952. “Pointing model procedure is giving higher RMS values than before (>1 rms, before was <1.0).” Responsable J. Urrutia. March 30,2009. It seems that stars used for the pointing procedure are not detected with enough precision and this could lead higher RMS. A complete 64-stars PoM was performed on April. Results were not yet conclusive. This LPRS will keep open at least until a new PoM is run. LPRS closed, solution: no time for this, and no problem has been reported.. Ismo Kastine, September 17,2009. ------------------------------------------------------------------------------ AREAS OF CONCERN: ------------------------------------------------------------------------------ PROJECTS FOR NEXT PERIOD: ------------------------------------------------------------------------------ USAGE OF TECHNICAL NIGHTS: ------------------------------------------------------------------------------ USAGE OF CALIBRATION PLAN NIGHTS (e.g. RESULTS, TRENDS, HEALTH CHECKS): ------------------------------------------------------------------------------ ---oOo---

2009 Q3-1 -- EFOSC2

La Silla Observatory: Quarterly Report Instrument EFOSC2 Name: Francisco C‡ceres Period: Q3 2009 (July -- September 2009) Date: 28 September 2009 ------------------------------------------------------------------------------ STATUS: - Personnel: IOT: Instrument Scientists: Colin Snodgrass Instrument TIO: Francisco C‡ceres System Engineer: Jorge Vilaza Instrument Friend: Francisco C‡ceres - General: * Alignment problem with the 1.0" slit, problem that is probably related to flexure of the instrument * Elongated images caused by the telescope oscillations, adapter rotator hard point. * A problem related to the time required by the sky flat to obtain the ADU level required. * MOS Y axis problem due to a loose cables. * There is request for spectroscopy, a sky flat template. * There is problem with the labels of some filters, these are the Schott filters in the EFOSC cabinet. - Instrument Report After the telescope and adapter rotator maintenances there is no new occurrence of the jumps and elongated images, but is pending a relation between the adapter temperature and the instrument focus. The problem of the Slit alignment probably is related to the instrument flexure, there is a workaround to check the alignment. Now is in development a template for doing sky flat field for spectroscopy. * Trends of instrument health-check parameters * Open or assigned LPRS (ordered according to decreasing priority): LPRS-009980:Slit aligment problem (On Hold) LPRS-009814:filter r#786 gives distorted PSF (Assigned) LPRS-010002:seqSh miissing on astro@wg5off (Open) LPRS-010051:Spectroscopic sky flat field template required (Assigned) LPRS-010052:order-sorting filters mislabelled (Assigned) * Other modified LPRS during Q3: LPRS-009977:EFOSC elongated images (Closed) LPRS-009959:Dome flat template aborts with FIERA error (Closed) LPRS-009946:sky flats (Closed) LPRS-009878:Hartmann test template (Closed) LPRS-010038:MOS axis Y not moving. (Closed) LPRS-010049:TCS and EFOSC OS failed (Closed) ------------------------------------------------------------------------------ AREAS OF CONCERN: * Related to LPRS-009980, is necessary a standard procedure to ensure that the 1.0" slit is aligned. * Related to LPRS-009814, is necessary to replace the r#787 filter. * Related to LPRS-009977, a relation between temperature adapter and focus problem. ------------------------------------------------------------------------------ PROJECTS FOR NEXT PERIOD: * Related to LPRS-010051, there is a template for spectroscopy sky flat, but is necessary include Tyson algorithm. * Related to LPRS-010052, filters could be numbered as ESO standard. ------------------------------------------------------------------------------ ---oOo---

2009 Q3-1 -- SOFI

La Silla Observatory : Quarterly Report Instrument : SOFI Name : Manuel Pizarro Period : Q3 2009 (July – September 2009) Date : 03 October 2009 STATUS : Personnel : no IOT : Instruments Scientists : no System Engineer : Peter Sinclaire Instrument Friend : Manuel Pizarro General * Usage of Sofi during Q3 2009 Visitor Mode : 2 -7 July 16 – 17 Jul 28 – Jul 1 Aug 7 - 12 Aug 14 – 15 Aug 5 – 11 Sep 13 – 14 Sep Instrument Report LPRS-010009 Date 7/9/2009 Sofi detector noise The noise in the SOFI detector appears again. Totally intermittent the last 3 hours of observation Date 7/14/2009 New occurrences of the random noise It seems to be the same at past year (LPRS-008173). Date 8/2/2009 New occurrence of the random noise. Since last statement made by JVA (change card), the noise has not been detected or reported. This is my Q3 2009 report... Manuel Pizarro La Silla TIO